OpenStudy (anonymous):

Use algebraic manipulations to evaluate lim as x approaches infinity: f(x)=((x^2)+2x-1)/(3+3(x^2))

OpenStudy (anonymous):

\[(\frac{x ^{2} + 2x - 1 }{ 3 + 3x ^{2} }) (\frac{ \frac{ 1 }{ x ^{2} } }{ \frac{ 1 }{ x ^{2} } })\] \[\frac{ 1 + \frac{ 2 }{ x } - \frac{ 1 }{ x ^{2} }}{ \frac{ 3 }{ x ^{2} } + 3 }\] then determine lim as x approaches infinity since N/(infinity) is 0 \[\frac{ 1+ 0 -0 }{ 0+3 }\] so the answer is \[\frac{ 1 }{ 3 }\] try checking my answer too, i might have missed something

OpenStudy (anonymous):

Thank you! The answer is 1/3 I just didn't know how to evaluate infinity.

OpenStudy (anonymous):

just always divide first the equation by the variable having the highest exponent

OpenStudy (anonymous):

Do you mean in general or when I am talking about infinity? I used the properties of limits to find the limit in each part and then was going to do the math but wasn't sure what value infinity represented.

OpenStudy (anonymous):

only when you're talkin' 'bout infitity